xx GiRaFFE: An Open-Source General Relativistic Force-Free Electrodynamics Code

GiRaFFE: An Open-Source General Relativistic Force-Free Electrodynamics Code

Zachariah B.xEtienne <zachetie *at* gmail *dot* com> Mew-Bing Wan Maria C.xBabiuc Sean T.xMcWilliams Ashok Choudhary

November 19, 2018

Abstract

GiRaFFE solves the equations of General Relativistic Force-Free Electrodynamics (GRFFE) using a high-resolution shock capturing scheme. It is based on IllinoisGRMHD, which is a rewrite of the Illinois Numerical Relativity (ILNR) group’s GRMHD code.

GiRaFFE is particularly good at modeling black hole magnetospheres. Its conservative-to-primitive solver has also been modified to check the physicality of conservative variables prior to primitive inversion, and move them into the physical range if they become unphysical.

Currently GiRaFFE consists of

the Piecewise Parabolic Method (PPM) for reconstruction,

the Harten, Lax, van Leer (HLL/HLLE) approximate Riemann solver, and